OpenStudy (anonymous):

Please help Find the difference of 9√7 and √28

OpenStudy (anonymous):

Ok, in order to subtract (or add) radicals, we need the same value underneath the square roots. So, lets look at the square root of 28. Remember, if 28 has a factor that is a perfect square, we can square root it and put it in front. Do you know which factor of 28 is a perfect square?

OpenStudy (anonymous):

Um honestly no I'm not to sure

OpenStudy (anonymous):

is it 7?

OpenStudy (anonymous):

Ok, we could try listing all the factors of 28: 1, 2, 4, 7, 14, 28 Which of those numbers is a perfect square? (Other than 1)

OpenStudy (anonymous):

7

OpenStudy (anonymous):

7 isn't a perfect square, because we can't take a whole number, square it, and get 7. Lets list some perfect squares: 2 squared is 4 <--- 3 squared is 9 4 squared is 16 4 is a perfect square because it is equal to 2 times 2

OpenStudy (anonymous):

okay so ?

OpenStudy (anonymous):

So, since the square root of 4 is 2, we can pull out the two to the front. Whats left in the radical is 28/4=7. So: \[\sqrt{28}=\sqrt{4} \times \sqrt{7}=2\sqrt{7}\]So what does this mean for the original problem? Well: \[9\sqrt{7} - \sqrt{28}=9\sqrt{7} - 2\sqrt{7} = ?\]

OpenStudy (anonymous):

7√7?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

yay !

OpenStudy (jdoe0001):

like roots add each other

OpenStudy (anonymous):

thank u